# A Bias-Correction Decentralized Stochastic Gradient Algorithm with Momentum Acceleration

### v1

@@ Abstract
Distributed stochastic optimization algorithms can handle large-scale data simultaneously and accelerate model training. However, the sparsity of distributed networks and the heterogeneity of data limit these advantages. This paper proposes a momentum-accelerated distributed stochastic gradient algorithm, referred to as Exact-Diffusion with Momentum (EDM), which can correct the bias caused by data heterogeneity and introduces the momentum method commonly used in deep learning to accelerate the convergence of the algorithm. We theoretically demonstrate that this algorithm converges to the neighborhood of the optimum sub-linearly irrelevant to data heterogeneity when applied to non-convex objective functions and linearly under the Polyak-Łojasiewicz condition (a weaker assumption than $\mu$-strongly convexity). Finally, we evaluate the performance of the proposed algorithm by simulation, comparing it with a range of existing decentralized optimization algorithms to demonstrate its effectiveness in addressing data heterogeneity and network sparsity.
@@ 1 Introduction
The distributed stochastic optimization (DSO) problem has attracted a great deal of attention in various domains such as the Internet of Things (IoT) (Ali et al., 2018), federated learning (Beltrán et al., 2023), and statistical machine learning (Boyd et al., 2011). Specifically, we consider the following problem
@@
where $F_{i}(\mathbf{x},\bm{\xi}_{i})$ is the loss function at agent $i$ given parameter $\mathbf{x}\in\mathbb{R}^{d}$, and $\bm{\xi}_{i}$ denotes the data on agent $i$, generated independent identically from the distribution $\mathcal{D}_{i}$. Each agent can exchange information with its neighbors, enabling collaborative decision-making to optimize the overall global objective function.
@@
A common way to solve such problems is the first-order method (Xin et al., 2020). Nedic and Ozdaglar (2009) proposed a sub-gradient optimization algorithm to solve the distributed optimization (DO), the deterministic counterpart of DSO. Leveraging the idea of stochastic approximation, these algorithms can be effectively adapted to address distributed stochastic optimization problems by estimating the gradient corresponding to the current parameter of each agent. Jiang et al. (2017) and Lian et al. (2017) propose a stochastic counterpart of Yuan et al. (2016), referred to as Distributed Stochastic Gradient Descent (DSGD), and studied their convergence properties on non-convex objective functions. Subsequently, a series of improved stochastic optimization algorithms are studied (Spiridonoff et al., 2020; Zhang and You, 2019; Tang et al., 2018).
@@
In a distributed scenario, the heterogeneity between each agent is an ineligible issue (Hsieh et al., 2020). Several works Yuan et al. (2021); Lin et al. (2021) pointed out that the DSGD can converge to a neighborhood of the optimal value, the radius of which is determined by the variance of the stochastic gradient and heterogeneity of the data for both strongly convex optimization and non-convex optimization. The network’s sparsity further exacerbates heterogeneity’s impact (Koloskova et al., 2020; Yuan et al., 2020). An effective method to address this issue is to modify the gradient update to approximate the average gradient. These algorithms can collectively be referred to as bias-correction algorithms. Zhang and You (2019) proposes distributed stochastic gradient tracking (DSGT), a widely used bias-correction algorithm, and proves its sub-linear convergence for non-convex objective functions. Its convergence property in the strongly convex case is established in Pu and Nedić (2021). In contrast to the DSGT, exact-diffusion (Yuan et al., 2020), also denoted as D2 in Tang et al. (2018), can eliminate the effect of data heterogeneity without aggregating extra information. In Alghunaim and Yuan (2022), a unified analysis of the bias-correction algorithms is conducted. The work Alghunaim and Yuan (2024) points out that the convergence rate of DSGT is slightly worse than ED/D2.
@@
To further enhance the convergence properties, a common strategy is to introduce the momentum method (Nesterov, 2013; Polyak, 1964). The works Yu et al. (2019); Gao and Huang (2020) introduce the idea of momentum into distributed stochastic optimization problems, and give the convergence rate in the non-convex case. However, heterogeneous bias cannot be eliminated through the introduction of momentum, making the convergence result less robust Yuan et al. (2021); Takezawa et al. (2023). Some methods (e.g. DecentLaM Yuan et al. (2021) and Quasi-Global Lin et al. (2021)) can compensate for this dilemma to some extent by modifying the momentum architecture. Introducing momentum into the bias-correction algorithm is a method that can fundamentally eliminate heterogeneity and retain the momentum acceleration effect. Several works have investigated how to incorporate momentum methods into DSGT and their convergence properties (Takezawa et al., 2023; Gao et al., 2023). Huang et al. (2024) introduce a loopless Chebyshev acceleration (LCA) technique to improve the consensus rate and prove its convergence properties under the PL condition case and general non-convex case. However, the momentum version of ED/D2, which serves as a bias correction algorithm, has been rarely studied, despite the theoretical results of Alghunaim and Yuan (2022) showing that ED/D2 should have better convergence properties in sparse networks and be more robust for highly heterogeneous data.
@@
Regarding the theoretical properties of the momentum-based algorithm, Liu et al. (2020) shows that the problem of centralized stochastic gradient descent with momentum possesses the same convergence bounds as stochastic gradient descent. Intuitively, momentum-based algorithms should have similar properties in distributed scenarios. However, existing theoretical analyses of momentum-based bias-correction algorithms pay little attention to this property and require additional constraints on step size or momentum selection, e.g. Takezawa et al. (2023); Gao et al. (2023) require step size $\alpha=\mathcal{O}((1-\lambda)^{2})$ and Huang et al. (2024) requires the momentum parameter $1-\beta=\mathcal{O}(1-\lambda)$. But Alghunaim and Yuan (2022) only requires a step size of $\alpha=\mathcal{O}(1-\lambda)$. Therefore, whether momentum-based algorithms achieve convergence rates comparable to those of the original algorithms is an important problem.
The contributions of this paper can be summarized as follows.
  1. 1.
    We propose a novel decentralized optimization algorithm incorporating momentum into the ED/D2 algorithm. This algorithm has the same structure as momentum SGD in the average sense and can accelerate the ED/D2 algorithm in the decentralized case. This ensures that our algorithm will not be affected by data heterogeneity and the network sparsity and will eventually converge to some neighborhood of the optimal solution, thus being more robust to the data heterogeneity. To the best of our knowledge, this work is the first to study the momentum version of ED/D2 and present a theoretical analysis.
  2. 2.
    This paper provides a novel technique to address the variance introduced by the complicated structure of stochastic gradients. Our approach establishes a tight upper bound on the convergence rate, which aligns with the results of the initial algorithm (ED/D2) in Alghunaim and Yuan (2022) for both non-convex and PL conditions. This alignment reinforces the conclusions drawn in Liu et al. (2020). Furthermore, the proof methodology can be extended to demonstrate the efficacy of other bias-correction algorithms.
@@
Table 1: General Non-Convex Convergence Rate. We focus on the condition when $1-\lambda$ is very small and treat our momentum parameter $\beta$ as a constant. Our conclusion of EDM eliminates the effect of the network sparsity, $\underline{\lambda}$, the minimum non-zero eigenvalue of $\mathbf{W}$, in Alghunaim and Yuan (2022). And when $\beta=0$, our conclusion is consistent with other work. Our conclusion explains that the ED/D2 can eliminate the influence of heterogeneity faster than the DSGT because the heterogeneous term (the bold part) of EDM is $(1-\lambda)^{-2}$, but DSGT is $(1-\lambda)^{-3}$.
@@ 2.1 Notations
Throughout this paper, we denote the parameter and data of agent $i$ at time $t$ as $\mathbf{x}_{i}^{(t)}\in\mathbb{R}^{d}$ and $\bm{\xi}_{i}^{(t)}\in\mathbb{R}^{p}$, respectively. Define $\mathbf{X}^{(t)}=(\mathbf{x}_{1}^{(t)},\cdots,\mathbf{x}_{n}^{(t)})^{\top}\in\mathbb{R}^{n\times d}$, $\bm{\Xi}^{(t)}=(\bm{\xi}_{1}^{(t)},\cdots,\bm{\xi}_{n}^{(t)})^{\top}$ For other variables, we use superscripts to represent parameters or data at time $t$ and subscripts to denote the labels of agents. We put $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$, $\mathbf{f}:\mathbb{R}^{n\times d}\to\mathbb{R}^{n}$, $\mathbf{f}(\mathbf{X})=(f_{1}(\mathbf{x}_{1}),\cdots,f_{n}(\mathbf{x}_{n}))^{\top}\in\mathbb{R}^{n}$. $F_{i}:\mathbb{R}^{d}\times\mathbb{R}^{p}\to\mathbb{R}$, $\mathbf{F}:\mathbb{R}^{n\times d}\to\mathbb{R}^{n}$, $\mathbf{F}(\mathbf{X},\bm{\Xi})=(F_{1}(\mathbf{x}_{1},\bm{\xi}_{1}),\cdots,F_{n}(\mathbf{x}_{n},\bm{\xi}_{n}))^{\top}$. The gradient of $\mathbf{F}$ w.r.t. $\mathbf{X}$ is defined as $\nabla\mathbf{F}(\mathbf{X},\bm{\Xi})=(\nabla F_{1}(\mathbf{x}_{1},\bm{\xi}_{1}),\cdots,\nabla F_{n}(\mathbf{x}_{n},\bm{\xi}_{n}))^{\top}$, and similarly for $\nabla\mathbf{f}(\mathbf{X})$. Let $\mathbf{1}_{d}=(1,\cdots,1)^{\top}\in\mathbb{R}^{d}$ denote the all-ones vector. The mean value of all node parameters is given by $\bar{\mathbf{x}}=\frac{1}{n}\mathbf{X}^{\top}\mathbf{1}_{n}\in\mathbb{R}^{d}$. Additionally, we define the matrix $\overline{\mathbf{X}}=\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\mathbf{X}\in\mathbb{R}^{n\times d}$. We define the average of loss functions with the local parameter as $\overline{\mathbf{F}}:\mathbb{R}^{n\times d}\times\mathbb{R}^{n\times p}\to\mathbb{R}$, given by $\overline{\mathbf{F}}(\mathbf{X},\bm{\Xi})=\frac{1}{n}\mathbf{F}(\mathbf{X},\bm{\Xi})^{\top}\mathbf{1}_{n}$. The mean of the individual loss functions $f_{i}(\mathbf{x}_{i})$ is denoted as $\bar{\mathbf{f}}:\mathbb{R}^{n\times d}\to\mathbb{R}$ with $\bar{\mathbf{f}}(\mathbf{X})=\frac{1}{n}\mathbf{f}(\mathbf{X})^{\top}\mathbf{1}_{n}$ The objective function $f:\mathbb{R}^{d}\to\mathbb{R}$ with $f(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\mathbf{x})$. For the momentum, let $\mathbf{m}_{i}^{(t)}$ be the momentum of agent $i$ at time $t$. Let $\mathbf{M}^{(t)}=(\mathbf{m}_{1}^{(t)},\cdots,\mathbf{m}_{n}^{(t)})^{\top}$ and $\bar{\mathbf{m}}^{(t)}=\frac{1}{n}(\mathbf{1}_{n}^{\top}\mathbf{M}^{(t)})^{\top}$. Denote the $\ell_{2}$-norm of $\mathbf{x}\in\mathbb{R}^{d}$ as $\|\mathbf{x}\|$, the Frobenius norm of matrix $\mathbf{A}$ as $\|\mathbf{A}\|_{\mathrm{F}}$ and the spectral norm as $\|\mathbf{A}\|_{\mathrm{op}}$.
@@ 2.2 Assumptions
In the decentralized setting, each agent in the network can only communicate with its neighbors to aggregate information. The network is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where the set of nodes is $\mathcal{V}=\{1,\cdots,n\}$, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ represents the connections between neighboring nodes, and we use $\mathcal{N}_{i}=\{j:(i,j)\in\mathcal{E}\}$ to denote the neighbor of the agent $i$.
@@
Let $\lambda=\|\mathbf{W}-\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}\|_{\mathrm{op}}$, which is the second largest eigenvalue of $\mathbf{W}$. From (1) and (2), we have $\lambda<1$ (Sayed et al., 2014). And $(1-\lambda)^{-1}$ is referred to as the spectral gap of the communication matrix $\mathbf{W}$ (Nedić et al., 2018; Huang et al., 2024), which is related to the efficiency of the consensus. The relationship between network structure and spectral gap is discussed in detail in Nedić et al. (2018). Take a ring graph for example, each agent is arranged in a ring and can only communicate with its two neighbors, in which case the spectral gap of the network will be $O(n^{2})$. In general, as the number of agents increases, the network tends to become sparser, which can lead to a larger spectral gap, subsequently hindering the convergence speed.
Note that (1) and (2) also imply that the smallest eigenvalue of $\mathbf{W}$ is greater than $-1$ (Sayed et al., 2014). Therefore, for any symmetric doubly stochastic matrix $\mathbf{W}$, Assumption 1 (3) can be satisfied through the following transformation: $\widetilde{\mathbf{W}}=\frac{\mathbf{W}+\mathbf{I}}{2}$.
@@
This assumption is widely used in distributed stochastic gradient descent algorithms (Alghunaim and Yuan, 2022; Huang et al., 2024), which mainly restricts the objective function to have a certain smoothness, to ensure that each step of gradient descent is relatively gentle.

### v2

@@ Abstract
Distributed stochastic optimization algorithms can process large-scale datasets simultaneously, significantly accelerating model training. However, their effectiveness is often hindered by the sparsity of distributed networks and data heterogeneity. In this paper, we propose a momentum-accelerated distributed stochastic gradient algorithm, termed Exact-Diffusion with Momentum (EDM), which mitigates the bias from data heterogeneity and incorporates momentum techniques commonly used in deep learning to enhance convergence rate. Our theoretical analysis demonstrates that the EDM algorithm converges sub-linearly to the neighborhood of the optimal solution, the radius of which is irrespective of data heterogeneity, when applied to non-convex objective functions; under the Polyak-Łojasiewicz condition, which is a weaker assumption than strong convexity, it converges linearly to the target region. Our analysis techniques employed to handle momentum in complex distributed parameter update structures yield a sufficiently tight convergence upper bound, offering a new perspective for the theoretical analysis of other momentum-based distributed algorithms.
@@ 1 Introduction
The distributed stochastic optimization (DSO) problem has garnered significant attention across various domains, including the Internet of Things (IoT) (Ali et al., 2018), federated learning (Beltrán et al., 2023), and statistical machine learning (Boyd et al., 2011). In this paper, we specifically address the following optimization problem:
@@
where $F_{i}(\mathbf{x},\bm{\xi}_{i})$ represents the loss function at agent $i$ given the parameter $\mathbf{x}\in\mathbb{R}^{d}$, and $\bm{\xi}_{i}$ denotes the data on agent $i$, which is independently and identically generated from the distribution $\mathcal{D}_{i}$. Each agent is capable of exchanging information with its neighbors, thus facilitating collaborative decision-making to optimize the overall global objective function.
@@
A common approach to solving DSO problems is through first-order methods (Xin et al., 2020). Notably, Nedic and Ozdaglar (2009) proposed a sub-gradient optimization algorithm for distributed optimization (DO), which serves as the deterministic counterpart of DSO. By leveraging the concept of stochastic approximation, these algorithms can be effectively adapted to tackle distributed stochastic optimization problems by estimating the gradient corresponding to the current parameters of each agent. Jiang et al. (2017) and Lian et al. (2017) introduced a stochastic variant of Yuan et al. (2016), referred to as Distributed Stochastic Gradient Descent (DSGD), and investigated its convergence properties for non-convex objective functions. Following this, a series of improved stochastic optimization algorithms were developed (Spiridonoff et al., 2020; Zhang and You, 2019; Tang et al., 2018).
@@
In a distributed scenario, data heterogeneity among agents presents a significant challenge (Hsieh et al., 2020). Several works Yuan et al. (2021); Lin et al. (2021) have indicated that DSGD can converge to a neighborhood of the optimal value, with the radius of this neighborhood determined by the variance of the stochastic gradient and the heterogeneity of the data for both strongly convex optimization and non-convex optimization. The sparsity of the network further exacerbates the impact of heterogeneity (Koloskova et al., 2020; Yuan et al., 2020). A viable strategy to mitigate this issue is to modify the gradient update to approximate the average gradient, collectively referred to as bias-correction algorithms. For instance, Zhang and You (2019) proposed Distributed Stochastic Gradient Tracking (DSGT), a widely used bias-correction algorithm, and demonstrated its sub-linear convergence for non-convex objective functions. Its convergence properties in the strongly convex case were established in Pu and Nedić (2021). In contrast, exact-diffusion (Yuan et al., 2020), also denoted as D2 in Tang et al. (2018), effectively eliminates the effects of data heterogeneity without requiring additional information aggregation. A unified analysis of the bias-correction algorithms is conducted in Alghunaim and Yuan (2022), and Alghunaim and Yuan (2024) noted that the convergence rate of DSGT is slightly inferior to ED/D2.
@@
To enhance convergence properties, a common strategy is to incorporate the momentum method (Nesterov, 2013; Polyak, 1964). Works by Yu et al. (2019) and Gao and Huang (2020) have introduced momentum into distributed stochastic optimization problems, providing convergence rates in the non-convex case. However, the introduction of momentum does not eliminate heterogeneous bias, which can render the convergence results less robust Yuan et al. (2021); Takezawa et al. (2023). Some methods, such as DecentLaM Yuan et al. (2021) and Quasi-Global Lin et al. (2021), partially address this issue by modifying the momentum architecture. Introducing momentum into the bias-correction algorithm presents a method that can fundamentally eliminate heterogeneity while retaining the benefits of momentum acceleration. Several studies have explored the incorporation of momentum methods into DSGT and their convergence properties (Takezawa et al., 2023; Gao et al., 2023). Huang et al. (2024) introduced a loopless Chebyshev acceleration (LCA) technique to improve the consensus rate and prove its convergence properties under the PL condition case and general non-convex case. Nonetheless, the momentum version of ED/D2, which functions as a bias-correction algorithm, has received limited attention, despite the theoretical results of Alghunaim and Yuan (2022) indicating that ED/D2 should exhibit superior convergence properties in sparse networks and greater robustness in the presence of highly heterogeneous data.
@@
Regarding the theoretical properties of the momentum-based algorithms, Liu et al. (2020) demonstrated that centralized stochastic gradient descent with momentum possesses the same convergence bounds as standard stochastic gradient descent. However, existing theoretical analyses of momentum-based bias-correction algorithms often overlook this property and impose additional constraints on step size or momentum selection. For instance, Takezawa et al. (2023); Gao et al. (2023) require step size $\alpha=\mathcal{O}((1-\lambda)^{2})$ and Huang et al. (2024) require the momentum parameter $1-\beta=\mathcal{O}(1-\lambda)$. In contrast, Alghunaim and Yuan (2022) only requires a step size of $\alpha=\mathcal{O}(1-\lambda)$. Therefore, whether momentum-based algorithms can achieve convergence rates comparable to those of the original algorithms remains an important question.
@@
Table 1: General Non-Convex Convergence Rate. We focus on the condition when $1-\lambda$ is very small and treat the momentum parameter $\beta$ as a constant. Our conclusion regarding EDM eliminates the impact of the network sparsity $\underline{\lambda}$, the minimum non-zero eigenvalue of $\mathbf{W}$, as stated in Alghunaim and Yuan (2022). Additionally, when $\beta=0$, our conclusion aligns with Yuan et al. (2020). Our conclusion demonstrates that the ED/D2 can eliminate the influence of heterogeneity more rapidly than the DSGT, since the heterogeneous term (the bold part) of EDM is scales with $(1-\lambda)^{-2}$, whereas DSGT scales with $(1-\lambda)^{-3}$.
@@
The contributions of this paper can be summarized as follows.
  1. 1.
    We propose a novel decentralized optimization algorithm that incorporates momentum into the ED/D2 framework. This algorithm can accelerate the ED/D2 algorithm in decentralized scenarios. And also ensures that our approach will not be affected by data heterogeneity and the network sparsity and will eventually converge to some neighborhood of the optimal solution, thus being more robust to the data heterogeneity. To the best of our knowledge, this work is the first work to investigate the momentum version of ED/D2 and to provide a theoretical analysis. Table 1 shows the theoretical advantages of our algorithm.
  2. 2.
    This paper introduces a new technique to address the variance introduced by the complicated structure of stochastic gradients. Our method establishes a tight upper bound on the convergence rate, which aligns with the findings of the original ED/D2 algorithm as presented in Alghunaim and Yuan (2022), applicable to both non-convex and PL conditions. This alignment reinforces the conclusions drawn by Liu et al. (2020). Furthermore, our proof methodology can be extended to demonstrate the effectiveness of other momentum-based bias-correction algorithms.
@@ 2.1 Notations
Throughout this paper, we denote the parameter and data of agent $i$ at time $t$ as $\mathbf{x}_{i}^{(t)}\in\mathbb{R}^{d}$ and $\bm{\xi}_{i}^{(t)}\in\mathbb{R}^{p}$, respectively. Define $\mathbf{X}^{(t)}=(\mathbf{x}_{1}^{(t)},\cdots,\mathbf{x}_{n}^{(t)})^{\top}\in\mathbb{R}^{n\times d}$ and $\bm{\Xi}^{(t)}=(\bm{\xi}_{1}^{(t)},\cdots,\bm{\xi}_{n}^{(t)})^{\top}$. For other variables, we use superscripts to indicate parameters or data at time $t$ and subscripts to denote the labels of agents. We put $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$, $\mathbf{f}:\mathbb{R}^{n\times d}\to\mathbb{R}^{n}$, $\mathbf{f}(\mathbf{X})=(f_{1}(\mathbf{x}_{1}),\cdots,f_{n}(\mathbf{x}_{n}))^{\top}\in\mathbb{R}^{n}$. $F_{i}:\mathbb{R}^{d}\times\mathbb{R}^{p}\to\mathbb{R}$, $\mathbf{F}:\mathbb{R}^{n\times d}\to\mathbb{R}^{n}$, $\mathbf{F}(\mathbf{X},\bm{\Xi})=(F_{1}(\mathbf{x}_{1},\bm{\xi}_{1}),\cdots,F_{n}(\mathbf{x}_{n},\bm{\xi}_{n}))^{\top}$. The gradient of $\mathbf{F}$ w.r.t. $\mathbf{X}$ is defined as $\nabla\mathbf{F}(\mathbf{X},\bm{\Xi})=(\nabla F_{1}(\mathbf{x}_{1},\bm{\xi}_{1}),\cdots,\nabla F_{n}(\mathbf{x}_{n},\bm{\xi}_{n}))^{\top}$, and similarly for $\nabla\mathbf{f}(\mathbf{X})$. Let $\mathbf{1}_{d}=(1,\cdots,1)^{\top}\in\mathbb{R}^{d}$ denote the all-ones vector. The mean value of all node parameters is given by $\bar{\mathbf{x}}=\frac{1}{n}\mathbf{X}^{\top}\mathbf{1}_{n}\in\mathbb{R}^{d}$. Additionally, we define the matrix $\overline{\mathbf{X}}=\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\mathbf{X}\in\mathbb{R}^{n\times d}$. We define the average of loss functions with the local parameter as $\overline{\mathbf{F}}:\mathbb{R}^{n\times d}\times\mathbb{R}^{n\times p}\to\mathbb{R}$, given by $\overline{\mathbf{F}}(\mathbf{X},\bm{\Xi})=\frac{1}{n}\mathbf{F}(\mathbf{X},\bm{\Xi})^{\top}\mathbf{1}_{n}$. The mean of the individual loss functions $f_{i}(\mathbf{x}_{i})$ is denoted as $\bar{\mathbf{f}}:\mathbb{R}^{n\times d}\to\mathbb{R}$ with $\bar{\mathbf{f}}(\mathbf{X})=\frac{1}{n}\mathbf{f}(\mathbf{X})^{\top}\mathbf{1}_{n}$ The objective function $f:\mathbb{R}^{d}\to\mathbb{R}$ with $f(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\mathbf{x})$. For the momentum, let $\mathbf{m}_{i}^{(t)}$ be the momentum of agent $i$ at time $t$. Let $\mathbf{M}^{(t)}=(\mathbf{m}_{1}^{(t)},\cdots,\mathbf{m}_{n}^{(t)})^{\top}$ and $\bar{\mathbf{m}}^{(t)}=\frac{1}{n}(\mathbf{1}_{n}^{\top}\mathbf{M}^{(t)})^{\top}$. Denote the $\ell_{2}$-norm of $\mathbf{x}\in\mathbb{R}^{d}$ as $\|\mathbf{x}\|$, the Frobenius norm of matrix $\mathbf{A}$ as $\|\mathbf{A}\|_{\mathrm{F}}$ and the spectral norm as $\|\mathbf{A}\|_{\mathrm{op}}$.
@@ 2.2 Assumptions
In the decentralized setting, each agent in the network can only communicate with its neighbors to aggregate information. The network is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where the set of nodes is $\mathcal{V}=\{1,\cdots,n\}$, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ represents the connections between neighboring nodes. We denote the neighbors of agent $i$ as $\mathcal{N}_{i}=\{j:(i,j)\in\mathcal{E}\}$.
@@
Let $\lambda=\|\mathbf{W}-\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}\|_{\mathrm{op}}$, which is the second largest eigenvalue of $\mathbf{W}$. From (1) and (2) in Assumption 1, we have $\lambda<1$ (Sayed et al., 2014). The quantity $(1-\lambda)^{-1}$ is referred to as the spectral gap of the communication matrix $\mathbf{W}$ (Nedić et al., 2018; Huang et al., 2024), which is related to the efficiency of the consensus among the agents. The relationship between network structure and spectral gap is discussed in detail in Nedić et al. (2018). For example, in a ring graph where each agent is arranged in a circle and can only communicate with its two neighbors, the spectral gap of the network will be $O(n^{2})$. Generally, as the number of agents increases, the network tends to become sparser, which can lead to a larger spectral gap and subsequently hinder the convergence speed.
It is worth noting that (1) and (2) in Assumption 1 also imply that the smallest eigenvalue of $\mathbf{W}$ is greater than $-1$ (Sayed et al., 2014). Therefore, for any symmetric doubly stochastic matrix $\mathbf{W}$, Assumption 1 (3) can be satisfied through the transformation $\widetilde{\mathbf{W}}=\frac{\mathbf{W}+\mathbf{I}}{2}$.
@@
This assumption is widely utilized in distributed stochastic gradient descent algorithms (Alghunaim and Yuan, 2022; Huang et al., 2024), which mainly restricts the objective function to have a certain smoothness, to ensure that each step of gradient descent is relatively gentle.